OpenStudy (anonymous):

If we toss two dice 1,000 times. How many times do we expect to have the sum of the two dice equal to 4? A. About 250 times B. About 167 times C. About 83 times D. About 42 times

OpenStudy (anonymous):

ways of getting sum of 4: (1,3), (2,2)

OpenStudy (hoblos):

when tossing two dice and add we have 36 possible outcome (6*6) we have 3 ways of getting sum of 4 (1,3) (2,2) (3,1) so the probability of getting a sum of 4 in each toss is 3/36 = 1/12 for 1000 tosses we expect 1000*1/12 = 83.33 times of having a sum of 4 83.33 is rounded to 83